OpenStudy (anonymous):

If a line contains the point (0, -1) and has a slope of 2, then which of the following points also lies on the line? (2, 1) (1, 1) (0, 1)

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@sarah_12

OpenStudy (precal):

do you know how to create a line using a slope and a point?

OpenStudy (anonymous):

no

OpenStudy (precal):

you are given the slope of 2 y=mx+b so your line is y=2x+b because slope is the "m" value

OpenStudy (precal):

(0,-1) happens to be your y intercept and that is your b value y=2x-1 that is your equation now all you have to do is figure out which of these points are on this line

OpenStudy (precal):

sub each point in one at a time for example is (2,3) on the line x=2 y=5 y=2x-1 5=2(2)-1 5=4-1 5=3 see how both sides are not equal, then it is false and thus (2,3) is not on the line

OpenStudy (precal):

test the three given points for your solution in the same fashion recall (x,y) First number is your x value and 2nd number is your y value

OpenStudy (anonymous):

2,1?

OpenStudy (anonymous):

no 1/1

OpenStudy (precal):

seems like you are guessing

OpenStudy (anonymous):

no ik what it is becaus e you said 0,-1 is my intercept so 1 up 1 would be ...

OpenStudy (precal):

but your slope is not 1 so going up 1 is incorrect your slope is 2 therefore you should be going up 2, right 1

OpenStudy (precal):

it is best for you just to sub the values into the equation

OpenStudy (precal):

(0,-1) you could start at this point and then go up 2 right 1 to determine other points but then to find points below (0,-1) then you go in the opposite direction which is down 2 left 1

OpenStudy (precal):

have you found the point yet?

OpenStudy (anonymous):

yeah 2,1

OpenStudy (precal):

(2,1) x=2 y=1 y=2x-1 1=2(2)-1 1=4-1 1=3 no because 1 does not equal 3

OpenStudy (precal):

(1,1) x=1 y=1 y=2x-1 1=2(1)-1 1=2-1 1=1 this is the point you want